Find the range, the standard deviation, and the variance for the given samples. Round non-integer results to the nearest tenth.

2, 8, 9, 12, 16, 19, 21


1
Expert's answer
2022-04-22T05:39:56-0400

Range is 21-2=19

"\mu=(2+8+9+12+16+19+21)/7=12.4"

"D=\frac{\sum (x-\mu)^2}{n-1}=\frac{108.8+19.36+11.6+0.16+12.96+43.6+73.96}{6}=45.1"

"\sigma=\sqrt{D}=6.7"